OpenStudy (anonymous):

what is 6.4 as a fraction

OpenStudy (anonymous):

\[\frac{ 6.4*10 }{ 1*10 }\] \[\frac{ 64 }{ 10 }\] \[2(\frac{ 64 }{ 10 })\] \[\frac{ 32 }{ 5 }\] http://www.mathsisfun.com/converting-decimals-fractions.html